OpenStudy (anonymous):

The pH of a 0.25 M aqueous solution of hydrofluoric acid, HF, at 25.0 °C is 2.03. What is the value of Ka for HF? 6.0 × 10-5 3.5 × 10-4 1.1 × 10-9 2.0 × 10-9 none of the above

OpenStudy (anonymous):

\[pH = pK_a-\log_{10}[HF]\]plug in the numbers

OpenStudy (anonymous):

\[2.03=pK_a-\log_{10}(0.25)\\ pK_a=? \]

OpenStudy (anonymous):

2.28?

OpenStudy (anonymous):

OH wait! I think I did that wrong.

OpenStudy (anonymous):

Okay, so i'm taking pH then adding the log of (.25) correct? When you put log10 is that different from the log button on my calc that just says log?

OpenStudy (anonymous):

log = \(\log_{10}\)

OpenStudy (anonymous):

you should get pKa = 1.428

OpenStudy (anonymous):

gotcha, so the answer i get when i add 2.03 and log(.25) is 1.4279 for pKa

OpenStudy (anonymous):

then I have to change that to Ka by inverse log, right?

OpenStudy (anonymous):

Nevermind, thanks tho. I guessed and got it right.

OpenStudy (anonymous):

\[ pH=-\log_{10}[H^+]\\ [H^+]=10^{-2.03}=0.00933\\ [F^-]=[H^+]=0.00933\\ [HF]=0.25-0.00933\\ \;\\ K_a=\frac{[H^+][F^-]}{[HF]}\\ K_a=\frac{0.00933\times0.00933}{0.25-0.00933}\\ \Large\boxed{K_a=3.6\times10^{-4}} \]